A beginning tennis class has twice as many students as the advanced class. The intermediate class has three more students than the advanced class. How many students are in the advanced class if the total enrollment for the three tennis classes is 43?
Answer by oberobic(2304)

a = advanced class size
b = beginner class size
i = intermediate class size
.
Total students is 43.
a + i + b = 43
.
Beginning class is twice the size of the advanced class.
b = 2a
.
Intermediate class has 3 more students than the advanced class.
i = a+3
.
substitute
.
a +a+3 +2a = 43
4a = 40
a = 10
.
b = 2a = 20
.
i = a+3 = 13
.
check
10+20+13=43
.
Answer: There are 10 students in the advanced class.
.
Done.